Operation Summary for October 19, 2009 - October 21, 2009 |
Monday Oct. 19 The day shift began with Tevatron (TeV) personnel and operators investigating the TeV quench, with Operations monitoring the Recycler stash (106.3E10), with the Antiproton Source stacking (21.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store. During the day shift, TeV personnel continued to investigate the loss of store 7278. Operations reported that a TeV quadrupole power supply (D0Q2) developed an internal LCW leak; EE Support personnel repaired the leak. Operations began the TeV turn-on and checked the orbits. Operations put the TeV in HEP shot setup at 11 AM and established store 7282 at 2:30 PM with an average initial luminosity (AIL) of 172.8E30. During the evening shift, a Booster RF station (BRF18), a Linac Klystron RF station (KRF3), and a MI sweeper magnet power supply tripped off; they all reset. Tuesday Oct. 20 The midnight shift began with Operations monitoring store 7282 (66.56E30) and the stash (242.2E10), with Pbar stacking (18.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7282 at 6:30 AM. Operations put the TeV into shot setup, but then reported a short delay to swap out the cryo system wet engine controller cards at TeV sectorA1. During the day shift, Operations established store 7283 at 9:15 AM with an AIL of 269E30. Various beam studies were conducted throughout the shift. Operations received permission to deliver beam to the Meson M01 enclosure. During the evening shift, the Linac controls front end (LIP611) suffered a hardware failure that required a new crate and power supply. Wednesday Oct. 21 The midnight shift began with Operations monitoring store 7283 (71.18E30) and the stash (324.8E10), with Pbar stacking (25.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7283 at 2 AM. The heat exchanger for the D0 #2 helium compressor was back flushed. Operations put the TeV into shot setup and established store 7284 at 4:15 AM with an AIL of 251E30. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow Preaccelerator personnel to flush the I- Source column. |
|
More Information Comments and Suggestions |
| last modified 10/21/2009 email Fermilab |
| Security, Privacy, Legal | |